OpenStudy (anonymous):

simplify 6 divided the sqrt of 14

OpenStudy (aravindg):

sqrt 14=sqrt 7*2

OpenStudy (aravindg):

6=3*2

OpenStudy (anonymous):

\[6/\sqrt{14}?\]

OpenStudy (anonymous):

yes @mustry

OpenStudy (anonymous):

rationalize denominator

OpenStudy (anonymous):



OpenStudy (anonymous):

\[6\sqrt{14}/14\]

OpenStudy (anonymous):

\[3\sqrt{14}/7\]

OpenStudy (anonymous):